OpenStudy (anonymous):

how do you work these out What is the mass of a cannonball if a force of 2,500 N gives the cannonball an acceleration of 200 m/s2? A. 12.5 kg B. 15.0 kg C. 20 kg D. 25 kg

OpenStudy (anonymous):

can anyone help me out?

OpenStudy (hoblos):

Newton's second law says \[\sum F = m \times a\] here we have only one force F=2500 N and a= 200 m/s2 can you calculate m ?

OpenStudy (anonymous):

um

OpenStudy (anonymous):

i don't know how but what i think it might be 15

OpenStudy (anonymous):

because what i did is 2 times 15 = to 25 even though it it 2500

OpenStudy (hoblos):

we have F = m*a thus m = F/a replace f by2500 and a by 200 what do you get ?

OpenStudy (anonymous):

2500 =m*200? thus m=2500/200=15?

OpenStudy (hoblos):

it's not 15, recalculate it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

12.5

OpenStudy (hoblos):

tell me what you get to check it for you

OpenStudy (hoblos):

yeah now it's correct

OpenStudy (anonymous):